OpenStudy (anonymous):

What is the derivative of 2^3x?

OpenStudy (cwrw238):

y = 2^(3x) = 8^x ln y = ln 8^x = x ln 8 1/y * dy/dx = ln 8 dydx = y ln 8 = 8^x ln 8